12) An 85 g arrow is fired from a bow whose string exerts an average force of 105 N on the arrow over a distance of 75 cm. What is the speed of the arrow as it leaves the bow? 8 A) 40.0 m/s B) 43.0 m/s C) 45.5 m/s D) 50.2 m/s E) 39.4 m/s
Answer #1

Using the Newton's law of motion

F=ma. Here m=85g=85×10^(-3)kg, F=105N

so acceleration a=F/m

a=105/85×10^(-3)

a=1.2353×10^3 m/s^2=1235.3 m/s^2

using equation of the motion

  V^2=u^2+2ax here u=0, a=1235.3 m/s^2 ,

x=75cm=0.75 m. So

V^2 =0+2×1235.3×0.75

V^2=1852.95

V=43.05 m/s

correct Answer is (B).